OpenStudy (anonymous):

Part 1: Create a difference of squares binomial or a perfect square trinomial that can be factored multiple times. (1 point) Part 2: Provide the factors of this polynomial. (2 points) Part 3: Explain, in complete sentences, the process you used to create the binomial/trinomial. (2 points) I don't really understand this at all. but I made up (x+3)(x+3)

OpenStudy (anonymous):

well a difference of squares would look like \[a ^{2} - b^{2} = (a - b)(a + b)\] so maybe change it to (x - 3)(x + 3) = x^2 - 9 the factors are simply (x - 3) and (x + 3)